A Very Lotharingien Paper:

A TENTATIVE ANSWER to the CASTEL-SAN PIETRO PROBLEM

by Bagarret KREWERLIENCO (Univ. of Adelaidelberg) and Peter Paul Roman BARNABHOFER (Univ. of Königfeld)

Summary. The authors give a simple but unuseful algorithm for quasi g-circulant embedding of skew trideterminants in the theory of patroids over chaotic sets, with applications to the breakfast-lunch-dinner problem.

Acknowledgments: The authors wish to thank:

prof. Andrè Numbers (Penntransylvania St. Univ.) and prof. Jacob Polynomials (Summer School of Cocumbinatorics)

for suggestions, straightenings and decompositions. They wish to thank also:

for Tableaux: Young & Yang Inc.

for Animals: the Zoo “G. Viennot”, Bordeaux. for various numbers: Euler & Genocchi ltd.

Editor’s Note: This paper is not large enough to contain statements, proofs and mistakes of the paper. Interested people can ask authors, or directly Mr. Lothaire, if they do not exist.